[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Sep 19, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Mon Sep 19 16:12:54 UTC 2016


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. TAFB Forecaster Gladys Rubio could not "Publish to Official" on AWIPS  
LX1 or LX2. She called NCF to report the problem and the issue was assigned  
Trouble Ticket #203. In the meantime, she sent out forecast from the Backup  
desk. NCF called back at 5:40 pm Sunday to inform TAFB that the problem was  
solved.

2. Mexico data outage (paraphrased from SDM log). At 0944Z TAFB Lead  
Forecaster Scott Stripling noticed a few missing synoptic obs over Mexico,  
(failures may have started at 06Z) and reported this to the SDM. Moments  
later, OMB/SS created a Trouble Ticket and Mexico station SMMX01 MXBA  
flagged red alert in Big Brother (CPBB). Tech Control sent a separate  
message to to the distant-end Point of Contact (Ms. Norma Angelica Tepoz  
Ortega) requesting support/assistance. At 1050Z the flag cleared to green  
alert on CPBB. Tech Control confirmed that the CP database showed an NWSTG  
receipt of product SMMX01 MXBA at 1045Z 9/18/16.
HSU:
1. At 5 pm EDT on Sunday, Sept. 18, the Hurricane Specialists sent the East  
Pacific Tropical Weather Outlook (TWOEP) multiple times. It posted to  
AWIPS, the GTWO updated and it printed, but it didn't show up on NHC's web  
site. TSB Web Developer Dave Zelinsky diagnosed the problem as a failure of  
the upstream NOAA Port LDM ingest, which missed both TWOs, probably because  
the NOAA Port feed was down entirely at the time the product was sent. To  
prevent future failures, he switched to NHC's new sockets feed from the  
Telecommunications Operations Center (TOC). HSU confirmed that this solved  
the problem.

2. WPC performed a back up test of HSU this morning at 15Z. The SDM (Senior  
Duty Meteorologist) called HSU to request that the guidance be resubmitted  
to the gyre (instead of the tide) supercomputing cluster. ATCF Developer  
Monica Bozeman modified the ${ATCFSCRIPTS}/nhc_putaids.sh shell script to  
reactivate ${SUPERCOMPUTERB} (gyre.ncep.noaa.gov) and followed up with a  
call to the SDM. All operational guidance (except the preliminary WSP and  
the preTCM) seems to have run correctly and posted to NHC's web site.
Other items:
1. Late yesterday afternoon, TAFB reported that the air conditioner had  
failed. HSU reported that air conditioning was restored at 11:30 pm.
NHC RFCs:
Submitted by NHC:

1. RFC #192 - 'text.pl' v3.0.1 - Remove code calls to generate blank  
placeholders for _experimental_ TAFB Satellite Rainfall forecast graphics  
on the NHC experimental webpage. [Note: the operational text products will  
not be impacted.] Was to be implemented on August 15 at 1500Z; delayed  
until the change was approved by the NHC experimental review board. That  
approval has been granted, so this is now scheduled for implementation on  
Tuesday, September 20 at 1500Z (11 am EDT). This RFC might be delayed  
again, depending on further review.

2. RFC #200 - nhc_awips.sh vatcf-nhc-2016-R1 - The ATCF KSH [Korn shell]  
script that formats text products and sends to AWIPS via PDS was failing to  
properly flag 'corrections'. Specifically: the ${ATCFSCRIPTS}/nhc_awips.sh  
script was looking for the 'correction code' (CCA, CCB, etc.) in the wrong  
location on the wrong header line. The code is being updated to properly  
extract correction code from product header line; also, header parsing  
logic will be revised generally to make the code more robust. To be  
implemented on Monday, September 19 at 1500Z (11 am EDT).

3. RFC #201 - create_pretcm_rcl.pl vatcf-nhc-2016-R1 - The "Pre" TCM/RCL  
text products, which are generated by the  
${ATCFSCRIPTS}/create_pretcm_rcl.pl Perl script, are not archived locally.  
The code is being revised to scp/save copies of the preTCM and RCL products  
to snook2 for internal NHC use. To be implemented on Monday, September 19  
at 1600Z (Noon EDT).

4. RFC #202 - Migrate C-shell scripts that plot significant wave height  
time series from various models at buoy locations. GIFs are posted to the  
NHC intranet server (snook). These scripts are being migrated from the  
pinfish server to the mirrored compute servers. This transition will  
facilitate troubleshooting and make the generation of these time series  
plots more robust/less prone to outages. To be implemented on Tuesday,  
September 20 at 1600Z (Noon EDT).

Related to NHC operations:

1. RFC #2323 - Change the DNS record for nhc.noaa.gov to support automatic  
redirect from nhc.noaa.gov to www.nhc.noaa.gov. To be implemented on  
Monday, September 19 at 1300Z (9 am EDT).

2. RFC #2329 - ens_tracker v1.1.9 - Ensemble tropical cyclone tracker and  
genesis jobs utilizing UKmet deterministic forecast data will be added to  
the ens_tracker package. At this time, the output tracks will only be sent  
to NHC's local directory on WCOSS. To be implemented on Tuesday, September  
20 at 1330Z (9:30 am EDT).

3. RFC #2330 - genesis_tracker v3.4.4 - The FNMOC tracker job is being  
added to the genesis_tracker family to extract tracks from FENS twice a  
day. The FNMOC deterministic model NAVGEM data location has been moved to  
/dcom/us007003. The tracker needs to be modified accordingly. To be  
implemented on Tuesday, September 20 at 1400Z (10 am EDT).

4. RFC #2240 [delayed] - product processor v1.1.0 - The production product  
parser for forecast.weather.gov is currently ignoring products that are  
corrections. This is a bug fix so that WWA (Watches/Warning/Advisory)  
corrections (COR) will show up on forecast.weather.gov. Was to be  
implemented on August 23 at 1400Z; postponed to September 7 at 1400Z due to  
CWD; postponed again to Monday, September 26 at 1400Z (10 am EDT) due to  
developer availability.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160919/4ef86855/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list